OpenStudy (anonymous):

Multiply. (4-√10)(2+5√6) Simplify your answer as much as possible.

OpenStudy (anonymous):

foil this actually

OpenStudy (dayakar):

\[(4-\sqrt{10}) (2+5\sqrt{6}) =4*2 +4*5\sqrt{6}-\sqrt{10}*2-\sqrt{10}*5\sqrt{6}\]

OpenStudy (dayakar):

\[=8+20\sqrt{6}-2\sqrt{10}-5*\sqrt{5}*\sqrt{2}*\sqrt{2}*\sqrt{3}\]

OpenStudy (dayakar):

\[=8+20\sqrt{6}-2\sqrt{10}-10\sqrt{15}\] here\[\sqrt{2}*\sqrt{2} =2\]

OpenStudy (anonymous):

Can you point to any khan academy / anywhere else that can teach me more about solving these types of problems? Thank you.

zepdrix (zepdrix):

Here is a simpler type problem from Khan Academy, https://www.youtube.com/watch?v=yAH3722GrP8 It involves a "variable", x, which your problem did not, but maybe it can still be of some help. @fishstix